Dublin University Rifle Club

Trinity College Dublin

DURC Budget Requirements

Mark Dennehy

May 20, 2001

1 Match Secretary Budget Proposal

1.1 Introduction

Given the experience gained from last year’s competitions, I have prepared the following proposed budget for DURC’s competition year from October 2000 to October 2001. This document should be reviewed by the incoming treasurer and a preliminary budget for the club for this year should be drawn up on the basis of the input from this document and similar documents from the equipment officer and other committee members.

1.2 Internal Competitions

1.2.1 Competitions

There are three internal competitions in the club :

• Air Rifle Competition

• Free Rifle Competition

• Sporter Rifle Competition

All three competitions have the same prizes :

Internal Competitions

First Place  G800G Medal w/gold trim £1.50
G309 Medal box £1.75 £3.25
Second Place G800S Medal w/silver trim £1.50
G309 Medal box £1.75 £3.25
Third Place G800B Medal w/bronze trim £1.50
G309 Medal box £1.75 £3.25
Total £9.75

1.2.2 Ladders There are three internal ladders in the club :

• Air Rifle Ladder

• Free Rifle Ladder

• Sporter Rifle Ladder All three will have the same prizes, but because we want to increase the interest in the

ladders, the prizes will be somewhat more substanstial :

Internal Ladders

First Place  291D Shield £9.20
Second Place 291B Shield £6.50
Third Place 291A Shield £5.45
Total £21.15

1.2.3 RO Knockout This is a new competition to try to encourage RO currency in all disciplines. It will take place over Hilary term, and all disciplines will be shot. Entry is mandatory for all Range Officers. One card to be shot per week, in a nominated discipline at any point in the week. The prize to be awarded at the AGM.

RO Knockout

First Place  H195 Pewter Hip Flash  £26.95
Second Place H814 Paperweight  £10.50
GB1 Base  £2.95 £13.45
Third Place H068B Silver Tray  £7.85
Total £48.25

1.2.4 Rapid Fire Competition

The rapid fire is a somewhat informal competition, but is also potentially the most popular if advertised properly. Hence, the prizes need to be somewhat better than for the internal competitions. However, we don’t have a huge budget, so the first prize is better and the second and third place prizes are standard.

Rapid Fire

First Place  352A Plaque  £7.95
Second Place G800S Medal w/silver trim £1.50
G309 Medal box £1.75 £3.25
Third Place G800B Medal w/bronze trim £1.50
G309 Medal box £1.75 £3.25
Total £14.45

1.2.5 Turkey Shoot

The Turkey shoot is the traditional christmas novelty match for the members of the club, as opposed to the Christmas Invitational, which is for external shooters and the top shooters in DURC. The event is more of a social one than a serious competition and so the prizes are the same as for the Rapid Fire competition. However, the cost of the competition is primarily the mulled wine and party snacks that are served during the competition. We can save a reasonable amount by capitalising on cooking skills in the club, but we can still expect to spend between 20 and 40 pounds. Exact figures have to wait until closer to the date.

Turkey Shoot

First Place  352A Plaque  £7.95
Second Place G800S Medal w/silver trim £1.50
G309 Medal box £1.75 £3.25
Third Place G800B Medal w/bronze trim £1.50
G309 Medal box £1.75 £3.25
Total £14.45

1.2.6 Grudge Match

The Grudge Match is usually held with one other club in the college, often during Trinity term and as it is supposed to be a social event, there are no prizes. Some money should be reserved to buy a round of drinks for the teams after the match though.

1.3 External Competitions

1.3.1 Postal Competitions

• UCESSA Free Rifle Postal, Indoor $28. 00

• UCESSA Free Rifle Postal, Outdoor

• UCESSA Air Rifle Postal $21. 00

• NTSA/ELEY Postal

1.4 Special Events

These are events which DURC hosts and which are open to the Irish shooting community in general. As such, they represent the major influence in DURC’s public image and need to go smoothly and have reasonable prizes to attract competitiors.

1.4.1 DURC 10m Air Open

Last year’s Air Open was de-registered because of the haphazard way it was run. This year it is re-registered, but we can’t really afford for it not to run smoothly. We should also have a perpetual trophy/shield for the open, if there isn’t one allready. Note that the cost of the perpetual trophy should be written off over the course of 11 years, i.e. approximately $10 per annum, including an allowance for the engraving of the current winner’s name. To avoid having to haul target boxes up the stairs in Luce Hall, and to discretely protest both last year’s fiasco and the current plans for the new range, we’re hiring UCD’s range for the event. Total cost for the hire is to be set at $300.

10m Air Open

Championship First Place  H848 1 Pint Tankard £3.60
GB2 Base £3.50  £7.10
343B Annual Shield  £43.75
Second Place  H847 \frac{1}{2} Pint Tankard £2.75
GB1 Base £2.95  £5.70
Third Place  H815 Shot Glass £3.45
GB1 Base £2.95  £6.40
Class A  First Place  316C Plaque  £5.25
Second Place  316B Plaque  £4.25
Third Place  316A Plaque  £3.20
Class B  First Place  295C Plaque  £5.25
Second Place  295B Plaque  £4.25
Third Place  295A Plaque  £3.20
Class C  First Place  E117B Plaque  £3.25
Second Place  E116B Plaque  £2.80
Third Place  E115B Plaque  £2.40
Class D  First Place  G800G Medal w/Gold trim £1.50
G309 Medal box £1.75  £3.25
Second Place G800S Medal w/Silver trim £1.50
G309 Medal box £1.75  £3.25
Third Place G800B Medal w/Bronze trim £1.50
G309 Medal box £1.75  £3.25
Range Rental  £300.00
Biscuits &c.  £20.00
Total £426.55

1.4.2 DURC Free Rifle 50m Open

Last year’s Open was resonably successful, though the barbecue which runs at the same time ran well over budget. Given what we learned from this, some changes should be made to the menu for the barbecue. Less salad, more hamburgers, hot dogs and kebabs. Potato salad is about all the salad that is required. Further, we need to have several people designated chefs on the day so that no-one gets sunstroke. Also, as for the Air open, we should have a perpetual shield. The one specified is very similar to that specified for the Air open, but different in color. Also, we should make sure that the date for the open and the date for the National squad training do not conflict this year as they did last year to prevent us losing half our attendees. Further, we will advertise the event a bit more this year. The menu for the barbecue is set up for 30-40 people, but this should be a reasonable estimate if we can get enought people interested. Rathdrum, Comber, East Antrim, Fassaroe and DRC should be encouraged to attend. The food budget for the barbecue should be about $100-120. The note on amortizing the cost of the perpetual shield from the Air open applies here also.

DURC 50m Free Rifle Open

Championship First Place  H848 1 Pint Tankard £3.60
GB2 Base £3.50  £7.10
344B Annual Shield  £43.75
Second Place  H847 \frac{1}{2} Pint Tankard £2.75
GB1 Base £2.95  £5.10
Third Place  H815 Shot Glass £3.45
GB1 Base £2.95  £6.40
Class A  First Place  316C Plaque  £5.25
Second Place  316B Plaque  £4.25
Third Place  316A Plaque  £3.20
Class B  First Place  295C Plaque  £5.25
Second Place  295B Plaque  £4.25
Third Place  295A Plaque  £3.20
Class C  First Place  E117B Plaque  £3.25
Second Place  E116B Plaque  £2.80
Third Place  E115B Plaque  £2.40
Class D  First Place  G800G Medal, Target Rifle Center, Gold trim £1.50
G309 Medal box £1.75  £3.25
Second Place G800S Medal, Target Rifle Center, Silver trim £1.50
G309 Medal box £1.75  £3.25
Third Place G800B Medal, Target Rifle Center, Bronze trim £1.50
G309 Medal box £1.75  £3.25
Range Rental  £200
Catering  £120
Total £425.95

Menu for barbecue

Hamburgers  100 patties

Vegetarian patties

100 buns 50 cheese slices

Pickled gherkin slices (one jar)

Hot dogs  30 frankfurters

30 buns

Kebabs  18 chicken breasts

12 sweet peppers

Potato Salad  One bowl
Coleslaw  Small bowl
Corn on the cob  20 half-cobs

Cob holders

Garlic bread  3 baguettes
Tortilla chips  2 large bags

Hot chilli dip Guacamole dip Cheese dip

Yogurt & Cucumber dip

Sweets  Bowl of pick’n’mix

Marshmallows for hot chocolate Ice cream

Drinks  Coke,Lemonade,Orange

Coffee, Tea Hot Chocolate

Condiments  Hamburger relish

Barbecue sauce Ketchup Mustard

1.4.3 DURC/UCDRC Colours Match

The colours is a match with a great deal of tradition attached, though little in the way of prizes. Perpetual trophies exist and do not need to be budgeted for, with the exception of engraving costs. However, there are other expenses that must be anticipated. Range hire, transport, food and drink must be allowed for, and coordinated with UCDRC, who share equally in the expenses. Further, a suitable gift should be purchased for Geoff as a token of appreciation. Previous years have seen large amounts of money spent due to lack of prior planning : This year purchasing a gift ahead of time will both save money and allow time for engraving, etc. Also, flowers or similar tokens should be purchased for Geoff’s family. Money must be set aside for the traditional round of drinks. Also, last year saw us giving medals to our colors squad in recognition of the time and effort devoted to training. This is to be encouraged and it was suggested by several people that we have medals made up for both UCD and DURC and present them at the end colors, cost to be divided with UCDRC.

Colors Match

Range Rental  £200
Gift for Geoff (Rock Tablet)  £41.85
Round of drinks  £100
G800G Medals w/gold trim x12 £18
G800S Medals w/silver trim x12  £18
G309 Medal Box x24 £42
Total £419.85

1.4.4 Trinity Top Ten 2001

The Top Ten is the regarded by the NTSA as being the pseudo-official postal competition. As such it has to be managed efficently. This does not entail a great deal of expense, except for the prizes. A perpetual trophy exists, but patches have to be made up for the prizes. Enquiries are being made with the original supplier of the Top Ten patches as to the cost of a small run of these. Cheap patches made from iron-on transfers will result in the competition not attracting any entries. Previous organisers have made a mess of this, and this year complaints have been made. This year’s prizes must be sorted out properly.

1.5 Supplier Choice

In previous years we have used Duiscé glass for prizes. This however, is currently beyond our budget and given the upcoming move to the new range it is likely to remain so, with the possible exception of the championship prizes for the two Open competitions. We have also used extremely cheap prizes last year when the budget was extremely limited. This however is counter-productive : no-one will participate in a competition where the prize is deemed worthless. However, several budget trophy retailers exist and last year we had good results with trophydirect.co.uk. We have also been promised a 15% discount and free engraving (except on silver orcrystal trophies) and free centres and activity figures on the trophies. I propose we use them as a sole supplier this year - however this will require that we purchase the prizes several weeks in advance. With planning this should not be an onerous task, and will allow for bulk-buying economies.

1.6 Final Budget

Prizes

G800 medals, various trims & engravings ×46 @ $1.50  $64.50
H195 Pewter Hip flask 1  $26.95
H814 Paperweight 1  $10.50
H068B Silver Tray 1  $7.85
352A Plaque ×2 @ $7.95  $15.90
H848 Tankard ×2 @ $3.60  $7.20
H847 Tankard ×2 @ $2.75  $5.50
H815 Shot Glass ×2 @ $3.45  $6.90
316[A,B,C] Plaque 6  $25.40
295[A,B,C] Plaque 6  $25.40
E11[5,6,7]B Plaque 6  $16.90
291[D,B,A] Shields 9  $63.45
343B Shield  $43.75
344B Shield  $43.75
Rock Tablet gift for Geoff  $41.85
GB1 Base 5 @ $2.95  $14.75
GB2 Base 2 @ $2.95  $6.90
G309 medal boxes ×46 @ $1.75  $80.50

$507.95

Less 15% Discount  $431.76

Range Rental

UCD DURC Air Open  $300
Wilkinstown Air Colours  $200
FSC 50m range training  $150
FSC DURC Free Open  $200

$850

Final Total $1281.76

Note that this does not include beverages or catering : and that there are still expenses that I have no details for as yet.

1.7 Intentions for Academic Year 2000/01

My intentions and goals for the upcoming year are as follows.

• Increase membership participation I intend to do this by giving out the schedule for competitions from the very beginning in the member’s handbook and on the web, including the prizes that are to be given out. Details will also be available on the web. Details of upcoming competitions will be posted on the email lists before the competitions : at one week beforehand, the day beforehand and at noon on the day of the competition. External competitions will be promoted similarly.

• Reintroduce competition squads Originally members qualified for external competitions on the basis of their average scores. Members above a certain level were inducted into a competition squad for Air, Free or Sporter rifle and had specific trainig periods, and were brought to the external competitions. Recently that idea has not been implemented. I intend to restart it this year however, with strict rules as to inclusion in the squad. Limiting the squad may seem counterproductive, but there are reasons. The arrangement will be to have squads of 12 for experienced shooters (more than one year in the club) and 6 for novices (first year shooting in that discipline). Knowing the maximum numbers that will be going to external competitions will ease the logistics of travel to these competitions and the notion of an elite group will challenge those in the club who are not merely plinking away casually, for whom the internal competitions will suffice. Again, the squad will be highly promoted from the member’s handbook onwards, and all members will wind up in photographs taken at events.

• Promote club events I intend to keep results, competition pages and news stories and articles up to date on the webpage, to keep announcements and reports flowing on the email lists and to write articles for both the student newspaper and the NTSA newsletters.

2 System Administrator / Website Administrator Budget Proposal/Report

The systems used in the club at present are adaquate to their tasks. This coming year will see a 10Mbps network installed on all three machines :

durc.tcd.ie This is the older Dell system which runs Linux.

scatt.tcd.ie This is the 486 which sits longside the Air rifle firing point and is used for the scatt machine.

playstation.tcd.ie This is the new machine which basically just plays games.

This 10Mbps network, running on coaxial cable, is allready installed and is comprised of obsolete CS Dept. cable and ISA network cards which were being thrown out but which are sufficent for our purposes. This will allow us to move the printer from playstation to durc and share it among all three machines. This will allow, for example, a shooter to print out a record of a SCATT training session which before would have been almost impossible.

durc.tcd.ie has been heavily upgraded. A new motherboard and CPU (now a Pentium 200MMX from a Pentium 90), two new network cards and an extra 1.2Gb hard disk have been installed. Its disks have been formatted and Linux installed. The second network card will be for the connection to the college network which will then be available to the other machines, as durc can act as a gateway. durc’s main function however, is to act as the club’s main server. It will run the webserver for the new website, which allows us to have more dynamic content such as bulletin boards, polls of users’ opinions and webcam broadcasts during competitions. It will run an email server and all lists will be moved to this machine. This will allow more lists past the current two, such as seperate announcement lists for Air, Free and Sporter rifle external competitions and so forth.

Further, with network access allowing the server to communicate to the outside, we can begin to move to a less labour-intensive management scheme for the club. Currently, all reports are paper-based and require duplicated effort on the part of the club’s officers. An online booking system for details, combined with databases on the club’s members will allow for the elimination of a large amount of effort and consequently a large amount of errors on the part of officers. Ammunition control in Front Arch, however, will remain paper-based both as a security precaution and as a check to this system. This system may seem optional for the club at its present level, however with the move to the new range in two or three years there will be an associated rise in membership coupled with a reduction in available time on the range and this will necessitate that a more efficent system be adopted. Rather than be caught at that point, it’s a better idea to develop the system now and refine it.

However this brings with it two main consequences. First is the financial outlay:

Network Connection  Initial Installation Fee  $200
Line rental (1 year)  $200
Domain name registration (optional)  $60
Image Capture (options) Quickcam  $60
WinTV  $60
Digital Camera  $150
Printer ink cartidges  $30
Final Total  $550 to 650

Secondly, this system will necessitate that the club from now on have an additional officer - that of sysadmin. Until this year that role has been a partial one, done on an informal basis. The management of the website and management system will require that this become a full officer’s position, and that that officer be chosen on the basis of his or her abilities. Obviously the sysadmin must understand, or be capable of quickly learning, how the system functions, and be able to update it as required. I am keeping written notes on the systems design : any future sysadmin must do the same.

A full report on the proposed management system is in preperation.

3 Miscalleanous Suggestions

In addition to the above budget proposals, I submit several suggestions :

• A fund to purchase christmas/birthday gifts for members of staff to maintain awareness of the Club on a personal level. A birthday card won’t break the bank - but sending one on the right day would show the relevant member of the sports department that we remembered. NTSA members should also be included on this list, as should Geoff (who always provides the Colours venue) and any coaches that we get to train DURC people.

• That Kathy and DUCAC be kept up-to-date by the treasurer to prevent the poor relations between DURC and DUCAC that the previous years have seen.

• The club needs to have an emergency ’war chest’ for coping with unanticipated expenses such as the Kurt Thune shooting jackets purchased last year. But this must take second place to setting the club up for the move to the new rifle range.

• The Department of Justice, Law Reform and Equality is conducting an internal review of the licencing arrangements for pistols. Should this go successfully, we should quickly move to re-certify the range for pistols and purchase some novice model air and smallbore pistols.